OpenStudy (anonymous):

Prove: tan2 θ•cos2 θ + cos2 θ = 1.

geerky42 (geerky42):

I would factor \(\cos^2\theta\) out.

geerky42 (geerky42):

And remember, \(\sin^2\theta +\cos^2\theta = 1\), so dividing both sides by \(\cos^2\theta\) will get you \(\underline{\tan^2\theta + 1 = \sec^2\theta}\). This ought to help you.

OpenStudy (anonymous):

is that answer? LOL :)

geerky42 (geerky42):

\(\tan^2\theta\cos^2\theta + \cos^2\theta = \cos^2\theta (\tan^2\theta + 1) = \cdots ~?\)

OpenStudy (anonymous):

is it 1